efeideue, soued 6. Assume that as of August 1, 3,000 units of flat panel displays have been produced and sold during the currCrystal Displays Inc. recently began production of a new product, flat panel displays, which required the investment of $1,306. A. Prepare a differential analysis of the proposed sale to Maple Leaf Visual Inc. Refer to the lists of Labels and AmountCan not figure out the variable manufacturing costs?
Answer #1

Variable manufacturing costs = (118+32+51)*900 = 180900

Reject order Accept Order Differential effect on income
Revenues 0 204300 204300
Costs:
Variable manufacturing costs 0 -180900 -180900
Income (Loss) per unit 0 26 26

Income per unit = (204300 -180900)/ 900 = $26 per unit